OpenStudy (hitaro9):

Let d be the distance from a point (x, y) on the line y = 11x − 8 to the point (0,0). Write d^2 as a function of x.

OpenStudy (hitaro9):

I also need to find the closest point on the line to 0,0.

OpenStudy (anonymous):

For the first question, the two points are (0,0) and (x, 11 x -8) Apply the distance formula to find \(d^2\)

OpenStudy (anonymous):

\[ d^2=(x-0)^2+ (11x -8 -0)^2=122 x^2-176 x+64 \]

OpenStudy (anonymous):

Minimize the last answer with respect to x

OpenStudy (anonymous):

\[ g(x) = d^2=122 x^2-176 x+64\\ g'(x)= 244x- 176\\ g'(x) =0 \implies x =\frac{44}{61} \implies y= 11 x -8 = -\frac 4{61} \]

OpenStudy (hitaro9):

Alright. Thank you. I'll see if that works. I appreciate the help.

OpenStudy (anonymous):

Did you understand it?

OpenStudy (hitaro9):

Yeah. I think so. You're using the formula for distance between two points right?

OpenStudy (hitaro9):

I haven't used that in forever. Hadn't thought of it.

OpenStudy (anonymous):

yes